freezing point depreesion

A 1.2436g sample of an unknown nonelectrolyte is dissolved in 10.9303g of benzophenone produces a solution that freezes at 43.2 degree C. If the pure benzophenonemelted at 48.1 degree C, what is the molecular weight of the unknown compound? The Kf for benzophenone is 9-8 degree C/m.
Answer #1

Step1 Let M be molar mass of solute ; Moles of solute =1.2463/M

Step2 ΔTf= 4.9 ; Kf= 9.8 C/m

Step3 Molality of solute =(1.2463/M)(1000/10.9303)

Step4 ΔTf = Kf xMolality ; 9.8x1.2463x1000/Mx 10.9303 =4.9

Step5 M= 228.04

answered by: Carter